Operations Report – February 17th

SOL: 7

Name of person filing report: Shannon Rupert (for Crew 189)

Non-nominal systems: Power System Control

Notes on non-nominal systems: Nothing new to report

Generator (hours run): Off at 8 am, on at evening (sorry no exact time).

Solar— 79 % SOC

Diesel – not checked %

Propane – not checked psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – not quite 10 gallons. Some were used to fill the tanks for ATV training.

Water (trailer) – 0 gallons.

Water (static) – 400 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: yes

Oil Added? no

ATV Fuel Used: not recorded Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: All four ARVs were used for training and practice around the campus. Atila took one of the 350’s up Cow Dung Road for samples.

Deimos rover used: Yes

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Sojourner rover used: Yes

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: Yes

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Opportunity rover used: Yes

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: No

HabCar used and why, where? Yes, to town to pick up food and meet crew

General notes and comments: The incoming crew did an awesome job of cleaning the Hab as soon as they arrived.

Summary of internet: Nothing to report
Summary of suits and radios: Nothing to report

Summary of Hab operations: The installation of the stairs, ladders, and railings is complete. While we were going to send photos tonight, the crew would like to wait until tomorrow to send good ones.

I took a shower in the Hab today as my hot water heater is currently broken. I was shocked by the horrible experience. There seemed to be no way to stop the water from getting too hot, then too cold, over and over. I wasted a lot of water just working with the water temperature. After my shower, I turned the hot water heater temperature down to what I think is a good temperature for using the
hot water only (much cooler than the scalding hot temperature that it was on). I will test tomorrow to make sure this is working. If not, we will have to find a solution or we will just waste a crazy amount
of water.

Summary of GreenHab operations: The hydroponics system had a leak and most of the water was lost overnight. We are currently troubleshooting a solution for the project.

Summary of ScienceDome operations: We did more lab work this morning and afternoon working on the halophile research. Eight samples x3 have been plated on 10% NaCl LB.

Summary of RAM operations: Atila will be working in the RAM beginning tomorrow.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing at this time